two particles of mass m each are moving in a horizontal circle with same angular speed .if both string are of same length then the ratio of tension in the string T1/T2=?
 

Grade:11

2 Answers

Arun
6 years ago
T2 equal to Mw2(2r)And on the inner mass particle we write T1-T2=Mw2r by balance the force on T1Now put the value of T2 T1-Mw2(2r)=Mw2rT1=3Mw2r Now T1/T2=3Mw2r/2Mw2r
Tejus
6 years ago
I am not getting the above solution if w are equal also length of string I.e. radius is equal then ratio should be1 to 1